### v3

# Spin polarized enantio-sensitive multipolar photoelectron currents

Abstract
Photoelectron circular dichroism (PECD) manifests as a forward-backward asymmetry of electron emission in the direction orthogonal to the light polarization plane via one-photon ionization of chiral molecules with circularly polarized light. Multi-polar ‘PECD’ currents, i.e., currents resolved along multiple directions, have also been predicted using two mutually-orthogonal linearly polarized light with carrier frequencies $\omega$ and $2\omega$. These currents arise from the interference between the one- and two-photon transitions. Here, we will show that photoelectron spin detection already reveals enantio-sensitive multi-polar currents in the one-photon regime since the two axes can be marked by the photoelectron momentum $\boldsymbol{\hat{k}}$ and spin-detection axis $\boldsymbol{\hat{s}}$. Specifically, we consider one-photon ionization of an isotropic ensemble of randomly oriented chiral molecules and show that the direction of the resulting photoelectron current is enantio-sensitively ‘locked’ to the photoelectron’s spin, which is mediated by two mechanisms. First, is the Bloch pseudovector which enables a collinear locking forming either a spin-sink or source for opposite enantiomers. Second, is the spin torque pseudovector that enables orthogonal locking forming a spin vortex in the polarization plane that rotates in opposite directions for opposite enantiomers. The former effect is a spin analog of photoelectron vortex dichroism (Phys. Rev. Lett. 129, 233201, 2022) wherein the detected photoelectron spin encodes molecular chirality while the latter is reminiscent of the Rashba effect in solids.
Photoelectron circular dichroism (PECD) heralded the “dipole revolution” in chiral discrimination: chiral discrimination without using chiral light [25, 5]. It was first predicted by Ritchie [31] and first detected by Böwering et. al. [6]. Nowadays, PECD is very well established theoretically, and also shown to yield strong enantiosensitive signals across several molecular species [29, 17, 37, 18, 21, 16, 19, 20, 22, 35, 11, 30, 36, 4]. In PECD, the photoionization of an isotropic ensemble of randomly oriented chiral molecules via circularly polarized light results to a forward-backward asymmetry in the net photoelectron current in the direction of light polarization. Fundamentally, it can be understood as a manifestation of a geometric magnetic field introduced in Ref. [25], and its emergence leads to new enantio-sensitive observables in photoionization [5, 23].
Geometric magnetism in photoionization [23] and photoexcitation [27, 32] of chiral molecules addresses the dynamical origin of enantiosensitive observables in one or multiphoton ionization. Its central object is the geometric propensity field:
$\boldsymbol{\vec{B}}_{\boldsymbol{\vec{k}}}\equiv i\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}}^{*}\times\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}},$ (1)
where, $\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}}$ is the photoelectron dipole field. The propensity field underlies several classes of such observables that originate from its multipole moments [23]: (i) the net propensity field $\boldsymbol{\vec{\Omega}}\equiv\int d\Theta_{k}\boldsymbol{\vec{B}}_{\boldsymbol{\vec{k}}}$; (ii) the net radial component of the propensity field $B^{\parallel}\equiv\int d\Theta_{k}(\boldsymbol{\hat{k}}\cdot\boldsymbol{\vec{B}}_{\boldsymbol{\vec{k}}})$; and (iii) the spherical multipole moments of the longitudinal $B_{l,m}^{\parallel}\equiv\int d\Theta_{k}(\boldsymbol{\hat{k}}\cdot\boldsymbol{\vec{B}}_{\boldsymbol{\vec{k}}})Y_{l,m}$ and transversal field components , $B_{l,m}^{\perp,1}\equiv\int d\Theta_{k}\boldsymbol{\vec{B}}_{\boldsymbol{\vec{k}}}\cdot\nabla_{k}Y_{l,m}$, and $B_{l,m}^{\perp,2}\equiv\int d\Theta_{k}\boldsymbol{\vec{B}}_{\boldsymbol{\vec{k}}}\cdot(\boldsymbol{\hat{k}}\times\nabla_{k})Y_{l,m}$.
PECD belongs to Class II observables [24], .i.e, $\boldsymbol{\vec{j}}_{\text{PECD}}\propto B^{\parallel}$, while Class I observables have been explored in Ref. [23]. It was shown that the net propensity field $\boldsymbol{\vec{\Omega}}$ controls the enantio-sensitive orientation of the cations - an effect called molecular orientation circular dichroism (MOCD). Class III observables have tensorial nature and represent multipolar photoelectron currents, i.e., currents resolved along multiple directions. The properties of such observables emerging in two photon ionization by two-color fields due to the interference between the one- and two-photon transitions have been explored theoretically in Refs. [13, 14, 26]. The same two-color set-up enabled detection of its possible analogue in multiphoton regime [33]. However, these effects ignore the photoelectron spin.
Spin detection opens an exciting opportunity to reveal enantio-sensitive multipolar currents arising already in the one-photon regime. Indeed, the two axes needed to resolve such currents can be marked by the photoelectron momentum $\boldsymbol{\hat{k}}$ and spin-detection axis $\boldsymbol{\hat{s}}$, see Fig. 1. The analysis of the interplay of chirality and spin in photoionization of chiral molecules was pioneered by Cherepkov [9, 10, 12]. He identified a kinematic picture of spin polarization, predicting spin- and enantio-sensitive effects in one photon ionization for circularly and linearly polarized light. First experiments detecting spin polarization in multiphoton ionization have just appeared [2]; simulations predicting spin-polarization in PECD have also become possible [3].
Here, we use our formalism on spin-resolved one-photon photoionization of chiral molecules [15] to explore multipolar photoelectron currents, and show that these currents are enabled by two complementary geometric mechanisms that arise solely from electric dipole interactions and the geometric properties of the photoionization dipoles in real space and spin space. The first mechanism is mediated by the Bloch pseudovector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$, and operates in fields of arbitrary polarization, including fully isotropic polarization. The second mechanism is the spin-resolved propensity field $\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}}$ which is a natural extension of Eq. (1) and is now a matrix in spin space. It only operates in fields with polarization planes fixed in the laboratory frame, e.g., circularly or elliptically polarized light.
We also use the same synthetic chiral argon system employed in Ref. [15], which are constructed by combining excited-state orbitals:
$|\psi_{m,\mu}^{\pm}\rangle_{p}=\dfrac{1}{\sqrt{2}}\left(|4p_{m},\mu\rangle\pm|4d_{m},\mu\rangle\right).$ (2)
$|\psi_{m,\mu}^{\pm}\rangle_{c}=\dfrac{1}{\sqrt{2}}\left(|4p_{m},\mu\rangle\pm i|4d_{m},\mu\rangle\right).$ (3)
These sates $|\psi_{m,\mu}^{\pm}\rangle_{p}$ and $|\psi_{m,\mu}^{\pm}\rangle_{c}$ are inspired by analogous chiral hydrogenic states [25]. Unlike hydrogen, the multielectron core potential in argon breaks inversion symmetry but the synthetic chirality is stabilized by electron correlations. The resulting photoelectron currents are equivalent to earlier predictions of Cherepkov [9, 10, 12], allowing us to identify the dynamical origins of spin polarization in photoionization. The direct comparison of our approaches will be addressed elsewhere.
Using perturbation theory, the full spinor valued electron wave-function at the end of the ionizing pulse can be written as
$|\psi\rangle=|\psi_{o}\rangle+\sum_{I,\mu^{M}}\int d\Theta_{k}^{M}\,c_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}|I\Psi_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}^{(-)}\rangle,$ (4a)
$c_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}=i\left(\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}^{L}\cdot\boldsymbol{\vec{E}}^{L}\right)$ (4b)
where, $|\psi_{o}\rangle$ is the ground state of the molecule, $|I\Psi_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}^{(-)}\rangle$ is the fully spin-coupled continuum state with momentum $\boldsymbol{\vec{k}}^{M}$ for the ionic channel $I$, $\mu=\pm\frac{1}{2}$ labels the photoelectron spin projection onto the molecular z-axis, $\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}^{L}=\langle\Psi_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}^{(-)}|\boldsymbol{\vec{d}}^{L}|\psi_{o}\rangle$ is the spin-resolved transition dipole, and $\boldsymbol{\vec{E}}^{L}$ describes the light field [^1] Superscripts $L$ and $M$ are used to denote quantities in the laboratory and molecular frames, respectively. Vectors in the molecular frame, $\boldsymbol{\vec{a}}^{M}$, are transformed into the laboratory frame using the relation $\boldsymbol{\vec{a}}^{L}=R_{\rho}\boldsymbol{\vec{a}}^{M}$, where $R_{\rho}$ is the Euler rotation matrix. Expressions without any superscript are to be understood as fully defined in the molecular frame.. For brevity, we will drop the index $I$ in the succeeding expressions.
Figure 1: Specification of coordinates in the laboratory frame. The light field propagates along $\boldsymbol{\hat{z}}$. The unit vector $\boldsymbol{\hat{\epsilon}}$ is the polarization direction for linearly polarized light, while $\boldsymbol{\hat{\Xi}}$ is the direction of photon spin for circularly polarized light. Upon ionization, the photoelectron is ejected in the direction of $\boldsymbol{\hat{k}}$ with its spin measured parallel to $\boldsymbol{\hat{s}}$.
The momentum and spin resolved photoionization rate $W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)$ for a given orientation $\rho$ is obtained by projecting the full wavefunction Eq. (4) onto the scattering and ionic states with energy $\mathcal{E}$, and projecting the spin onto the axis $\boldsymbol{\hat{s}}^{L}$:
$\hat{\mathsf{P}}_{\mathcal{E}}$= $\sum_{\mu,\nu}\int d\Theta_{k}|\Psi_{\boldsymbol{\vec{k}},\mu}^{(-)}\rangle(\hat{\mathsf{P}}_{\boldsymbol{\hat{s}}})_{\nu,\mu}\langle\Psi_{\boldsymbol{\vec{k}},\nu}^{(-)}|.$ (5)
Here, $\hat{\mathsf{P}}_{\boldsymbol{\hat{s}}}=(\mathbb{I}+\boldsymbol{\hat{s}}\cdot\boldsymbol{\hat{\sigma}})/2$ is a spin projection operator with respect to $\boldsymbol{\hat{s}}$, and $\boldsymbol{\hat{\sigma}}$ is the vector of Pauli spin matrices. Performing the necessary operations, we obtain
$W^{L}$ $(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)$
$=$ $\dfrac{1}{2}\sum_{I,\mu,\nu}\int d\rho\left(\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\mu}^{L*}\cdot\boldsymbol{\vec{E}}^{L*}\right)\left(\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\nu}^{L}\cdot\boldsymbol{\vec{E}}^{L}\right)$
$\times\left(\delta_{\mu,\nu}+\boldsymbol{\hat{s}}^{L}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{L}\right).$ (6)
The photoelectron current conditioned on the spin detection axis $\boldsymbol{\hat{s}}^{L}$ can now be calculated as follows:
$\boldsymbol{\vec{j}}^{L}(\boldsymbol{\hat{s}}^{L})=$ $\dfrac{\int d\rho\int d\Theta_{k}^{L}W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)\boldsymbol{\vec{k}}^{L}}{\int d\rho\int d\Theta_{s}^{L}\int d\Theta_{k}^{L}W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)}$ (7)
wherein, $\int d\rho$, $\int d\Theta_{k}^{L}$, and $\int d\Theta_{s}^{L}$ denotes averaging over all molecular orientations, photoelectron momentum, and spin detection axis, respectively. Full details of the succeeding calculations are shown in Appendix A.
Let us first consider randomly oriented chiral molecules under isotropic illumination by linearly polarized light, i.e., instead of having a single direction $\boldsymbol{\hat{\epsilon}}^{L}$ to characterize to polarization direction (see Fig. 1), we introduce
$\boldsymbol{\vec{E}}^{L}=$ $E_{\omega}^{L}\left(\sin\theta_{p}\cos\varphi_{p}\boldsymbol{\hat{x}}^{L}+\sin\theta_{p}\sin\varphi_{p}\boldsymbol{\hat{y}}^{L}+\cos\theta_{p}\boldsymbol{\hat{z}}^{L}\right).$ (8)
Subsituting Eq. (8) into (7) and averaging over all polarization directions $\int d\Theta_{p}$, the resulting photoelectron current is now
$\boldsymbol{\vec{j}}_{\text{iso}}^{L}=$ $\dfrac{1}{3S_{0}}\left(\dfrac{1}{k}\int d\boldsymbol{\vec{\Theta}}_{k}^{M}\cdot\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}\right)\boldsymbol{\hat{s}}^{L},$ (9)
where, $S_{0}$ is the total yield
$S_{0}=\int d\Theta_{k}\left(\left|\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},\frac{1}{2}}\right|^{2}+\left|\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},-\frac{1}{2}}\right|^{2}\right).$ (10)
Equation (9) shows that the current is proportional to the flux of the vector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}$ through the surface of the energy shell ($d\boldsymbol{\vec{\Theta}}_{k}^{M}=d\Theta_{k}^{M}\boldsymbol{\hat{k}}^{M}k^{2}$). Mathematically, $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ is a Bloch pseudovector:
[…]
$\tilde{\varrho}=\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},\mu}^{*}\cdot\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},\nu}\right),$ (11b)
describing spin orientation in the degenerate two-level system formed by spin-up and spin-down continuum states populated by photoionization, which incorporates both populations and coherences within this system [15]. Here, $\tilde{\varrho}$ is the reduced density matrix which emerges after performing a partial trace over the spatial continuum states and degenerate ionization channels of the full density matrix, and averaging over all molecular orientations.
The current $\boldsymbol{\vec{j}}_{\text{iso}}^{L}$, Eq. (9), shows that photoelectrons with opposite spins are correlated to opposite enantiomers, i.e.,
$\boldsymbol{\vec{j}}_{\text{iso}}^{L(S)}(\boldsymbol{\hat{s}}^{L})=\boldsymbol{\vec{j}}_{\text{iso}}^{L(R)}(-\boldsymbol{\hat{s}}^{L}),$ (12)
and that the current is collinearly ‘locked’ to the direction of the spin-detection axis $\boldsymbol{\hat{s}}$. Hence, $\boldsymbol{\vec{j}}_{\text{iso}}^{L(S)}(\boldsymbol{\hat{s}}^{L})$ serves as the spin analog of photoelectron vortex dichroism (PEVD) [28]. Since, PEVD uses linearly polarized light to generate photoelectrons with orbital angular momentum (OAM) wherein the OAM has opposite signs for opposite enantiomers, Eq. (9) now suggests that this collinearly ‘locked’ photoelectron current can be enhanced by using light fields with either a well-defined polarization direction or polarization plane.
Figure 2: Schematic of the resulting photoelectron currents for (a,b) linearly, and (c-f) circularly polarized light in the laboratory frame. The solid and dashed blue vectors denote the opposite directions of spin for opposite enantiomers. (a-d) The currents $\boldsymbol{\vec{j}}_{\text{lin}}$ and $\boldsymbol{\vec{j}}_{r}$ are enabled by the Bloch vector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ which results to a collinear locking of the photoelectron current and spin $\boldsymbol{\vec{j}}\parallel\boldsymbol{\hat{s}}$. (e-f) The currents $\boldsymbol{\vec{j}}_{\text{PECD}}$ and $\boldsymbol{\vec{j}}_{\tau}$ are enabled by the spin-resolved propensity field $\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}}$. The PECD current is not spin-sensitive, while $\boldsymbol{\vec{j}}_{\tau}$ presents a three-way orthogonal locking of the photoelectron current and spin with the photon spin.
Repeating the same process for linearly polarized light
$\boldsymbol{\vec{E}}^{L}=E_{\omega}^{L}\boldsymbol{\hat{\epsilon}}^{L},$ (13)
we obtain the photoelectron current
$\boldsymbol{\vec{j}}_{\text{lin}}^{L}=$ $\dfrac{1}{5S_{0}}\left[\dfrac{1}{k}\int d\boldsymbol{\vec{\Theta}}_{k}^{M}\cdot\left(2\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}+\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}\right)\right]\boldsymbol{\hat{s}}^{L}$
$-\dfrac{1}{5S_{0}}\left[\dfrac{1}{k}\int d\boldsymbol{\vec{\Theta}}_{k}^{M}\cdot\left(\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}+3\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}\right)\right]\left(\boldsymbol{\hat{s}}^{L}\cdot\boldsymbol{\hat{\epsilon}}^{L}\right)\boldsymbol{\hat{\epsilon}}^{L}$ (14a)
$\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}=$ $\text{Re}\left[\sum_{\mu,\nu}\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M*}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{M}\right)\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{M}\right].$ (14b)
The additional vector $\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}$ is the directional bias introduced by the well-defined direction of the light polarization. Hence, the photoelectron current now arises as the flux of an effective Bloch pseudovector $\boldsymbol{\vec{\mathsf{S}}}=a\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}+b\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}$ through the surface of the energy shell.
Figure 3: Comparison of the photoelectron currents enabled by the Bloch vector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ for different chiral states [Eqs. (2)-(3)] via random illumination, linearly (a,b) and circularly polarized light (c,d) with either orthogonal or collinear detection geometry [see Fig. 1]. The rapidly oscillating behavior at higher values of $k$ are due to the Fano resonances, leading up to the ionization threshold for the 3s electrons [34, 8]
The current $\boldsymbol{\vec{j}}_{\text{lin}}^{L}$, Eq. (14), can be measured using two possible schemes: (i) collinear $\boldsymbol{\hat{s}}\parallel\boldsymbol{\hat{\epsilon}}$ and (ii) orthogonal $\boldsymbol{\hat{s}}\perp\boldsymbol{\hat{\epsilon}}$ detection geometry, see Fig. 2(a,b), respectively. Enhancement or reduction of $\boldsymbol{\vec{j}}_{\text{lin}}^{L}$ with respect to $\boldsymbol{\vec{j}}^{L}_{\text{iso}}$ is then controlled by $\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}$:
[…]
A comparison of the currents $\boldsymbol{\vec{j}}_{\text{iso}}^{L}$ and $\boldsymbol{\vec{j}}_{\text{lin}}^{L}$ for the states $|\psi_{+1,\frac{1}{2}}^{\pm}\rangle_{c}$ and $|\psi_{-1,\frac{1}{2}}^{\pm}\rangle_{p}$ is shown in Fig. 3(a,b). It can be seen that enhancement of the current is indeed possible but not always ensured as it is sensitive both to the chiral state and detection geometry. Figure 3(a) demonstrates that the orthogonal detection geometry enhances the photoelectron current but is reduced for collinear detection geometry. Meanwhile, Fig. 3(b) shows that both detection geometries enhances the resulting photoelectron current with the collinear detection geometry now having a larger enhancement.
Now, consider circularly polarized light
$\boldsymbol{\vec{E}}^{L}=E_{\omega}^{L}\dfrac{(\boldsymbol{\hat{x}}^{L}+i\xi\boldsymbol{\hat{y}}^{L})}{\sqrt{2}}$ (16)
where, $\xi=\pm 1$ is a dichroic parameter characterizing the direction of rotation of the light polarization vector. The resulting current will have the following components:
$\boldsymbol{\vec{j}}_{\text{circ}}^{L}=\boldsymbol{\vec{j}}_{r}^{L}+\boldsymbol{\vec{j}}_{\text{PECD}}^{L}+\boldsymbol{\vec{j}}_{\tau}^{L}.$ (17)
The term $\boldsymbol{\vec{j}}_{r}^{L}$ is enabled by the Bloch vector $\boldsymbol{\vec{\mathcal{S}}}^{M}_{\boldsymbol{\vec{k}}^{M}}$ and has a similar form as that of Eq. (14):
$\boldsymbol{\vec{j}}_{r}^{L}=$ $\dfrac{1}{10S_{0}}\left[\dfrac{1}{k}\int d\boldsymbol{\vec{\Theta}}_{k}^{M}\cdot\left(3\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}+\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}\right)\right]\boldsymbol{\hat{s}}^{L}$
$+\dfrac{1}{10S_{0}}\left[\dfrac{1}{k}\int d\boldsymbol{\vec{\Theta}}_{k}^{M}\cdot\left(\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}-3\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}\right)\right](\boldsymbol{\hat{s}}^{L}\cdot\boldsymbol{\hat{\Xi}}^{L})\boldsymbol{\hat{\Xi}}^{L},$ (18)
which is enantio-sensitive but not dichroic. The collinear locking between the photoelectron current and spin is still present can be measured either via collinear or orthogonal detection geometry with respect to the photon spin direction $\boldsymbol{\hat{\Xi}}^{L}=(-i\boldsymbol{\vec{E}}^{L*}\times\boldsymbol{\vec{E}}^{L})/|\boldsymbol{\vec{E}}^{L}|^{2}=\xi\boldsymbol{\hat{z}}^{L}$, see Fig. 2(c,d). Moreover, the current $\boldsymbol{\vec{j}}_{r}^{L}$ is contained in the coefficients $B_{1}$ and $B_{2}$ introduced by Cherepkov [9, 10, 12]. Figure 3(c) shows that there is negligible effect to the current compared to $\boldsymbol{\vec{j}}_{\text{iso}}^{L}$. Meanwhile, Fig. 3(d) demonstrates enhancement of the current but it is almost half of $\boldsymbol{\vec{j}}_{\text{lin}}^{L}$.
The terms $\boldsymbol{\vec{j}}_{\text{PECD}}^{L}$ and $\boldsymbol{\vec{j}}_{\tau}^{L}$ are enabled by the spin-resolved propensity field:
$(\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}})_{\mu,\nu}=i\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},\mu}^{*}\times\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},\nu}$ (19)
which is now a matrix in spin space, and a natural extension of Eq. (1). Explicitly, we have
[…]
which shows that PECD is not spin sensitive since it takes the sum of the spin-up and down photoelectrons, and proportional to the flux of the propensity field through the surface of the energy shell ($d\boldsymbol{\vec{\Theta}}^{M}=d\Theta_{k}^{M}\boldsymbol{\hat{k}}^{M}k^{2}$). Equation (20) is equal to the PECD current we derived in [24], to the original expression derived by Ritchie [31], and to the coefficient $D$ introduced by Cherepkov [9, 10, 12].
Figure 4: (a,b) Comparison of the components of $\boldsymbol{\vec{j}}_{\text{circ}}$ for different chiral states [Eqs. (2)-(3)] with $\boldsymbol{\hat{s}}\parallel\boldsymbol{\hat{y}}$.
The current
$\boldsymbol{\vec{j}}_{\tau}^{L}=$ $\dfrac{1}{4N}\left(\dfrac{1}{k}\int d\boldsymbol{\vec{\Theta}}_{k}^{M}\cdot\boldsymbol{\vec{\tau}}^{M}\right)\left(\boldsymbol{\hat{s}}^{L}\times\boldsymbol{\hat{\Xi}}^{L}\right)$ (21a)
$\boldsymbol{\vec{\tau}}_{\boldsymbol{\vec{k}}^{M}}^{M}=$ $\text{Tr}(\boldsymbol{\hat{\sigma}}^{M}\times\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}^{M}}^{M}).$ (21b)
is confined in the polarization plane which indicates transversal spin-polarization. This current is maximal when the photoelectron spin detection axis $\boldsymbol{\hat{s}}^{L}$ is orthogonal to both the direction of the current and photon spin $\boldsymbol{\hat{\Xi}}^{L}$. Equation (21) can also be equivalently written as the flux of the spin torque vector $\boldsymbol{\vec{\tau}}_{\boldsymbol{\vec{k}}^{M}}^{M}$ through the surface of the energy shell. It is also equivalent to the coefficient $C$ of Cherepkov [9, 10, 12], in which $\boldsymbol{\vec{j}}_{\tau}^{L}\propto A^{1,1}_{1,-1}-A^{1,-1}_{1,1}$. The enantio-sensitive nature of $\boldsymbol{\vec{j}}_{\tau}^{L}$ arises from the molecular factor $(\boldsymbol{\vec{\tau}}^{M}\cdot\boldsymbol{\vec{k}}^{M})$ which is a pseudoscalar: opposite enantiomers produce opposite spin vortices, such that spin polarization direction depends on photoelectron momentum in a way reminiscent of Rashba effect in solids.
Alternatively, $\boldsymbol{\vec{j}}_{\tau}^{L}$ can be written in terms of the spherical multipole moments of the the spin-resolved propensity field as follows:
$\boldsymbol{\vec{j}}_{\tau}^{L}=\dfrac{1}{4S_{0}}$ $\left\{\int d\Theta_{k}^{M}\,\text{Tr}\left[\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}^{M}}^{M}\cdot\left(\boldsymbol{\vec{k}}^{M}\times\nabla_{k}\vec{\mathbb{Y}}_{\boldsymbol{\hat{k}}^{M}}^{M}\right)\right]\right\}$
$\left(\boldsymbol{\hat{s}}^{L}\times\boldsymbol{\hat{\Xi}}^{L}\right)$ (22a)
$\vec{\mathbb{Y}}_{\boldsymbol{\vec{k}}^{M}}^{M}=\sqrt{\dfrac{4\pi}{3}}\begin{bmatrix}Y_{1,0}(\boldsymbol{\hat{k}}^{M})&\sqrt{2}Y_{1,-1}(\boldsymbol{\hat{k}}^{M})\\ -\sqrt{2}Y_{1,1}(\boldsymbol{\hat{k}}^{M})&-Y_{1,0}(\boldsymbol{\hat{k}}^{M})\end{bmatrix}$ (22b)
This indicates that $\boldsymbol{\vec{j}}_{\tau}^{L}$ belongs to the Class III enantio-sensitive observables. Hence, to find spin-sensitive effects, we have to consider multipolar photoelectron currents, i.e. currents resolved along two orthogonal directions. Without spin detection, multipolar enantio-sensitive currents could be induced using orthogonally polarised two-color fields [13, 33, 26, 5], and recorded via interference between the two photoionization pathways involving the interference of one-photon and two-photon transitions.
The three components of $\boldsymbol{\vec{j}}_{\text{circ}}^{L}$ become mutually orthogonal when $\boldsymbol{\hat{s}}\perp\boldsymbol{\hat{\Xi}}$, hence, can be measured independently [see Fig. 2(d-f)]. A comparison of each term in $\boldsymbol{\vec{j}}_{\text{circ}}^{L}$ is shown in Fig. 4. The vortex current $\boldsymbol{\vec{j}}_{\tau}^{L}$ is negligible for the state $|\psi_{1,\frac{1}{2}}^{\pm}\rangle_{c}$ while the radial current $\boldsymbol{\vec{j}}_{r}^{L}$ is almost eight times than PECD. Meanwhile, $\boldsymbol{\vec{j}}_{r}^{L}$ is negligible for the state $|\psi_{-1,\frac{1}{2}}^{\pm}\rangle_{p}$ but the vortex current $\boldsymbol{\vec{j}}_{\tau}^{L}$ is almost twice the PECD. Thus, spin detection opens an exciting opportunity to detect enantio-sensitive multipolar currents already in the one-photon regime, because the orthogonal axes can be marked by the photoelectron momentum and its spin.
In summary, we have shown how measurements of the photo-electron spin can be used to detect spin polarized enantio-sensitive currents. These currents are enabled by two geometric mechanism: (i) the Bloch pseudovector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ which enables collinear locking of the photoelectron current with the photoelectron spin and operators for arbitrary light fields, even isotropic illumination, and (ii) the spin-resolved propensity field $\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}}$ which enables orthogonal locking and only operators for light fields with a fixed polarization plane. We have also presented several detection schemes to measure these currents.
Appendix A Derivation of photoelectron currents
The photoelectron current conditioned on the spin detection axis $\boldsymbol{\hat{s}}^{L}$ is calculated as follows:
$\boldsymbol{\vec{j}}^{L}(\boldsymbol{\hat{s}}^{L})=$ $\dfrac{\int d\rho\int d\Theta_{k}^{L}W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)\boldsymbol{\vec{k}}^{L}}{\int d\rho\int d\Theta_{s}^{L}\int d\Theta_{k}^{L}W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)}$ (23a)
$W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)=$ $\dfrac{1}{2}\sum_{\mu,\nu}\int d\rho\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{L*}\cdot\boldsymbol{\vec{E}}^{L*}\right)\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{L}\cdot\boldsymbol{\vec{E}}^{L}\right)\left(\delta_{\mu,\nu}+\boldsymbol{\hat{s}}^{L}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{L}\right)$
$=$ $\sum A^{\ell,m_{\ell}}_{\ell_{s},m_{s}}Y_{\ell,m_{\ell}}(\boldsymbol{\hat{k}}^{L})Y_{\ell_{s},m_{s}}(\boldsymbol{\hat{s}}^{L}),$ (23b)
[…]
The vectors that appear on the right-hand side of Eq. (23b) can be grouped into two sets: (i) vectors that are fixed in the molecular frame such as the dipole transition vectors $\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}^{M}$, photoelectron momentum $\boldsymbol{\vec{k}}^{M}$, and photoelectron spin quantization axis $\boldsymbol{\hat{\sigma}}_{\mu_{1}^{M},\mu_{2}^{M}}^{M}$, and (ii) vectors that are fixed in the laboratory frame such as spin detection axis $\boldsymbol{\hat{s}}^{L}$ and the electric field $\boldsymbol{\vec{E}}^{L}$. This will then allow us to use the technique in Ref. [1] in evaluating the orientation averaging $\int d\rho$ such that the resulting quantity can be expressed as $\sum_{ij}g_{i}M_{ij}f_{j}$, where, $g_{i}$ and $f_{i}$ are rotational invariants that are constructed from the two sets of vectors and $M_{ij}$ is the coupling between the two rotational invariants. For our purposes, the following vector identities will be relevant:
$\int d\rho(\boldsymbol{\vec{a}}^{L}\cdot\boldsymbol{\vec{u}}^{L})\boldsymbol{\vec{b}}^{L}=\dfrac{1}{3}(\boldsymbol{\vec{a}}^{M}\cdot\boldsymbol{\vec{b}}^{M})\boldsymbol{\vec{u}}^{L}$ (24)
$\int d\rho(\boldsymbol{\vec{a}}^{L}\cdot\boldsymbol{\vec{u}}^{L})(\boldsymbol{\vec{b}}^{L}\cdot\boldsymbol{\vec{v}}^{L})\boldsymbol{\vec{c}}^{L}=\dfrac{1}{6}[(\boldsymbol{\vec{a}}^{M}\times\boldsymbol{\vec{b}}^{M})\cdot\boldsymbol{\vec{c}}^{M}](\boldsymbol{\vec{u}}^{L}\times\boldsymbol{\vec{v}}^{L})$ (25)
$\int d\rho(\boldsymbol{\vec{a}}^{L}\cdot\boldsymbol{\vec{u}}^{L})(\boldsymbol{\vec{b}}^{L}\cdot\boldsymbol{\vec{v}}^{L})(\boldsymbol{\vec{c}}^{L}\cdot\boldsymbol{\vec{w}}^{L})\boldsymbol{\vec{d}}^{L}=\dfrac{1}{30}\begin{bmatrix}(\boldsymbol{\vec{a}}^{M}\cdot\boldsymbol{\vec{b}}^{M})(\boldsymbol{\vec{c}}^{M}\cdot\boldsymbol{\vec{d}}^{M})\\ (\boldsymbol{\vec{a}}^{M}\cdot\boldsymbol{\vec{c}}^{M})(\boldsymbol{\vec{b}}^{M}\cdot\boldsymbol{\vec{d}}^{M})\\ (\boldsymbol{\vec{a}}^{M}\cdot\boldsymbol{\vec{d}}^{M})(\boldsymbol{\vec{b}}^{M}\cdot\boldsymbol{\vec{c}}^{M})\end{bmatrix}^{T}\begin{bmatrix}4&-1&-1\\ -1&4&-1\\ -1&-1&4\end{bmatrix}\begin{bmatrix}(\boldsymbol{\vec{u}}^{L}\cdot\boldsymbol{\vec{v}}^{L})\boldsymbol{\vec{w}}^{L}\\ (\boldsymbol{\vec{u}}^{L}\cdot\boldsymbol{\vec{w}}^{L})\boldsymbol{\vec{v}}^{L}\\ (\boldsymbol{\vec{v}}^{L}\cdot\boldsymbol{\vec{w}}^{L})\boldsymbol{\vec{u}}^{L}\end{bmatrix}$ (26)
Using Eqs. (24)-(25), we can easily evaluate the normalization factor as follows:
$\int d\rho$ $\int d\Theta_{s}^{L}\int d\Theta_{k}^{L}W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)$
$=$ $\int d\Theta_{s}^{L}\int d\Theta_{k}^{M}\int d\rho W^{L}(\boldsymbol{\hat{k}}^{M},\boldsymbol{\hat{s}}^{M},\rho)$
$=$ $\dfrac{1}{2}\int d\Theta_{s}^{L}\int d\Theta_{k}^{M}\sum_{\mu,\nu}\int d\rho\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{L*}\cdot\boldsymbol{\vec{E}}^{L*}\right)\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{L}\cdot\boldsymbol{\vec{E}}^{L}\right)\left(\delta_{\mu,\nu}+\boldsymbol{\hat{s}}^{L}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{L}\right)$
$=$ $\dfrac{1}{6}\int d\Theta_{s}^{L}\int d\Theta_{k}^{M}\sum_{\mu,\nu}\left[\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M*}\cdot\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{M}\right)|\boldsymbol{\vec{E}}^{L}|^{2}\delta_{\mu,\nu}\right]$
$+\dfrac{1}{12}\left\{\int d\Theta_{k}^{M}\sum_{\mu,\nu}\left[\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{M}\cdot\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M*}\times\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{M}\right)\right]\right\}\left[\left(\int d\Theta_{s}^{L}\boldsymbol{\hat{s}}^{L}\right)\cdot\left(\boldsymbol{\vec{E}}^{L*}\times\boldsymbol{\vec{E}}^{L}\right)\right]$
$=$ $\dfrac{1}{6}\left(\sum_{\mu}\int d\Theta_{k}^{M}\left|\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M}\right|^{2}\right)|\boldsymbol{\vec{E}}^{L}|^{2}$ (27)
The second line follows from the definition of a rotated function $W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)=W^{M}(\boldsymbol{\hat{k}}^{M},\boldsymbol{\hat{s}}^{M},\rho)$ [7], and we have interchanged the order of integration to perform change of variable $\boldsymbol{\hat{k}}^{M}=R_{\rho}^{-1}\boldsymbol{\hat{k}}^{L}$ and again interchanged the order of integration since $\boldsymbol{\hat{k}}^{M}$ is now an integration variable independent of $\rho$. Meanwhile, the last line follows from the vanishing of $\int d\Theta_{s}^{L}\boldsymbol{\hat{s}}^{L}$.
Let us now consider the spin-conditioned photoelectron current under isotropic illumination:
$\boldsymbol{\vec{j}}_{\text{iso}}^{L}=\dfrac{\int d\Theta_{p}\int d\rho\int d\Theta_{k}^{L}W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)\boldsymbol{\vec{k}}^{L}}{\int d\Theta_{p}\int d\rho\int d\Theta_{s}^{L}\int d\Theta_{k}^{L}W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)},$ (28)
where, $\int d\Theta_{p}$ denotes averaging over all orientations of the light field:
$\boldsymbol{\vec{E}}^{L}_{p}=E_{\omega}^{L}\left(\sin\theta_{p}\cos\varphi_{p}\boldsymbol{\hat{x}}^{L}+\sin\theta_{p}\sin\varphi_{p}\boldsymbol{\hat{y}}^{L}+\cos\theta_{p}\boldsymbol{\hat{z}}^{L}\right)$ (29)
The relevant quantity to calculate is the the numerator of Eq. (28), i.e.,
$\int d\Theta_{k}^{M}$ $\int d\Theta_{p}\int d\rho W^{M}(\boldsymbol{\hat{k}}^{M},\boldsymbol{\hat{s}}^{M},\rho)\boldsymbol{\vec{k}}^{L}$
$=$ $\dfrac{1}{2}\sum_{\mu_{1}^{M},\mu_{2}^{M}}\int d\Theta_{k}^{M}\int d\Theta_{p}\int d\rho\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{L*}\cdot\boldsymbol{\vec{E}}_{p}^{L*}\right)\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{L}\cdot\boldsymbol{\vec{E}}_{p}^{L}\right)\left(\delta_{\mu,\nu}+\boldsymbol{\hat{s}}^{L}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}\right)\boldsymbol{\vec{k}}^{L}$ (30)
Using Eq. (25), the first term of Eq. (30) simplifies into
$\dfrac{1}{2}\sum_{\mu}$ $\int d\Theta_{k}^{M}\int d\Theta_{p}\int d\rho\left|\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{L}\cdot\boldsymbol{\vec{E}}_{p}^{L}\right|^{2}\boldsymbol{\vec{k}}^{L}$
$=$ $\dfrac{1}{12}\left\{\left[\sum_{\mu}\int d\Theta_{k}^{M}\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M*}\times\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M}\right)\right]\cdot\boldsymbol{\vec{k}}^{M}\right\}\left[\int d\Theta_{p}\left(\boldsymbol{\vec{E}}^{L*}_{p}\times\boldsymbol{\vec{E}}^{L}_{p}\right)\right]=0,$ (31)
which vanishes after averaging over all orientations of the field. Similarly, it follows from Eq. (26) that the second term of Eq. (28) is now
$\dfrac{1}{2}\sum_{\mu,\nu}\int d\Theta_{k}^{M}$ $\int d\Theta_{p}\int d\rho\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{L*}\cdot\boldsymbol{\vec{E}}_{p}^{L*}\right)\left(\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\nu}^{L}\cdot\boldsymbol{\vec{E}}_{p}^{L}\right)\left(\boldsymbol{\hat{s}}^{L}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{L}\right)\boldsymbol{\vec{k}}^{L}$
$=\dfrac{1}{60}$ $\begin{bmatrix}\sum\int d\Theta_{k}^{M}\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M*}\cdot\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{M}\right)\left(\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{M}\cdot\boldsymbol{\vec{k}}^{M}\right)\\ \sum\int d\Theta_{k}^{M}\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M*}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{M}\right)\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{M}\cdot\boldsymbol{\vec{k}}^{M}\right)\\ \sum\int d\Theta_{k}^{M}\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M*}\cdot\boldsymbol{\vec{k}}^{M}\right)\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{M}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{M}\right)\end{bmatrix}^{T}\begin{bmatrix}4&-1&-1\\ -1&4&-1\\ -1&-1&4\end{bmatrix}\begin{bmatrix}\int d\Theta_{p}|\boldsymbol{\vec{E}}_{p}^{L}|^{2}\boldsymbol{\hat{s}}^{L}\\ \int d\Theta_{p}(\boldsymbol{\vec{E}}_{p}^{L*}\cdot\boldsymbol{\hat{s}}^{L})\boldsymbol{\vec{E}}_{p}^{L}\\ \int d\Theta_{p}(\boldsymbol{\vec{E}}_{p}^{L}\cdot\boldsymbol{\hat{s}}^{L})\boldsymbol{\vec{E}}_{p}^{L*}\end{bmatrix}$
$=\dfrac{1}{60}$ $\begin{bmatrix}g_{1}\\ g_{2}\\ g_{2}^{*}\end{bmatrix}^{T}\begin{bmatrix}4&-1&-1\\ -1&4&-1\\ -1&-1&4\end{bmatrix}\begin{bmatrix}f_{1}\\ f_{2}\\ f_{2}^{*}\end{bmatrix}$
$=\dfrac{1}{30}$ $\left(2f_{1}-\text{Re}[f_{2}]\right)g_{1}-\dfrac{1}{30}\left(f_{1}-3\text{Re}[f_{2}]\right)\text{Re}[g_{2}]-\dfrac{1}{6}\text{Im}[f_{2}]\text{Im}[g_{2}]$ (32)
Averaging over all orientations of the field we get:
$\dfrac{1}{2}\sum_{\mu,\nu}\int d\Theta_{k}^{M}$ $\int d\Theta_{p}\int d\rho\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{L*}\cdot\boldsymbol{\vec{E}}_{p}^{L*}\right)\left(\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\nu}^{L}\cdot\boldsymbol{\vec{E}}_{p}^{L}\right)\left(\boldsymbol{\hat{s}}^{L}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{L}\right)\boldsymbol{\vec{k}}^{L}$
$=$ $\dfrac{1}{18}|\boldsymbol{\vec{E}}^{L}|^{2}\left\{\int d\Theta_{k}^{M}\left[\sum_{\mu,\nu}\left(\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\mu}^{M*}\cdot\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\nu}^{M}\right)\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{M}\right]\cdot\boldsymbol{\vec{k}}^{M}\right\}\boldsymbol{\hat{s}}^{L}$
$=$ $\dfrac{1}{18}|\boldsymbol{\vec{E}}^{L}|^{2}\left[\int d\Theta_{k}^{M}\left(\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}\cdot\boldsymbol{\vec{k}}^{M}\right)\right]\boldsymbol{\hat{s}}^{L}.$ (33)
Thus,
$\boldsymbol{\vec{j}}_{\text{iso}}^{L}=\dfrac{1}{3S_{0}}\left[\int d\Theta_{k}^{M}\left(\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}\cdot\boldsymbol{\vec{k}}^{M}\right)\right]\boldsymbol{\hat{s}}^{L}$ (34a)
$S_{0}=\int d\Theta_{k}\left(\left|\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},\frac{1}{2}}\right|^{2}+\left|\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},-\frac{1}{2}}\right|^{2}\right).$ (34b)
The same procedure is done for both linearly and circularly polarized light.