### v4

# arXiv:2505.23460v4[quant-ph] 26 Feb 2026

Abstract
Chiral molecular structures enable time-odd pseudotensor that controls correlations between photoelectron spin and molecular orientation in photoionization [13]. We show that chiral structures also support time-even correlations between the photoelectron spin and its momentum. Fundamentally, these correlations can only be revealed in conditioned measurements. We argue that conditioned measurements are the sole origin of a broad class of phenomena termed chirality-induced spin selectivity (CISS), because all these phenomena require a correlated detection of the spin and another variable, e.g., electron current. We consider one-photon ionization of an isotropic ensemble of randomly oriented chiral molecules and show that the direction of the resulting photoelectron current is enantio-sensitively ‘locked’ to the photoelectron spin even under isotropic illumination. We also identify molecular frame spin texture created by photoionization in which its flux through the energy shell is the molecular pseudoscalar that quantifies the underlying spin-current correlations. Next, we consider the effect of the photon spin on these correlations and find that chiral structure also leads to triple correlations between the photoelectron momentum, its spin and the photon spin. Our results identify two molecular pseudovectors underlying the full complexity of the multitude of spin-conditioned measurements of the photoelectron current
Chirality induced spin selectivity (CISS) usually refers to a spin-polarized charge transport through chiral molecules [29], but in a wider context it underscores a recently discovered functionality of chiral molecules to filter photoelectron spin [12]. CISS is surrounded with intrigue and debates on its origin and strength which stimulate detailed simulations, simple models and new experiments [12]. CISS can be initiated either by photoionization [16] or by applying a voltage across a chiral molecular device [36]. The latter usually involves a substrate coated with a thin film of chiral molecules and a spin-sensitive probe, such as a magnetic conductive probe atomic force microscope (mCP-AFM). The tip of the microscope is magnetized to resolve the current conditioned on spin. Flipping the direction of tip’s magnetisation, one detects the difference in the current through the device.
This example shows that CISS related phenomena are fundamentally linked to conditioned measurements. Such measurements reveal correlations between some vectorial molecular property $\mathbf{v}$ and an electron spin $\mathbf{s}$: $\langle\mathbf{v}\otimes\mathbf{s}\rangle$. Such pseudoscalar correlations can only be enabled by chiral structures and can have different strength and manifestations depending on the specific measurement. Starting at the premises of conditioned measurements, we consider the simplest possible scenario of chiral spin-filtering: photoionization of an ensemble of randomly oriented chiral molecules. Our aim is to identify the origin of such pseudoscalar correlations. The simple scenario that has not yet been quantified is the one removing all external biases such as molecular orientation or polarization of the light field.
Consider one photon ionization of randomly oriented chiral molecules under isotropic illumination. To detect a net current, one needs to measure photoelectrons in all possible directions. Since the measurement, the system, and the light are completely isotropic, there will be no net current in the system. However, if one introduces a spin detection axis, $\boldsymbol{\hat{s}}$, it breaks the symmetry and a net current along $\boldsymbol{\hat{s}}$ is symmetry allowed (see Fig. 1). The current conditioned on spin can only have a component along the spin detection axis. Since the current is a time-odd polar vector and spin is a time-odd axial vector, the projection of the current on the spin axis is a time-even pseudoscalar. Its pseudoscalar nature dictates that such current can only emerge in chiral molecules; as a time even quantity, it does not require breaking of the time-reversal symmetry of the Hamiltonian. Thus, the current is absent without spin conditioning and emerges as soon as the measurement is conditioned (or postselected) on spin.
To address the strength and nature of the these pseudoscalar correlations, we use our geometric formalism developed for spin-resolved one-photon photoionization of chiral molecules [13]. It allows us to identify and quantify the molecular pseudoscalrs underlying both time-even spin-current correlations as well as time-odd triple correlations between the spin of the photoelectron, the photoelectron current, and the photon spin.
We show that these correlations are enabled by two complementary geometric mechanisms that arise solely from electric dipole interactions and the geometric properties of the photoionization dipoles in real space and spin space. The first mechanism is mediated by the momentum-resolved Bloch pseudovector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ defined below. It operates in fields of arbitrary polarization, including fully isotropic polarization. The second, time-odd, mechanism is mediated by the spin-resolved propensity field $\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}}$ which is a natural extension of its spin-averaged precursor [25, 3] responsible for photoelectron circular dichroism [30, 27, 4, 21, 18] and is also defined below. This molecular property requires breaking time-reversal symmetry and can be activated by photon spin.
To quantify our results we use synthetic chiral argon system employed in Ref. [13], which are constructed by combining excited-state orbitals:
$|\psi_{m,\mu}^{\pm}\rangle_{p}=\dfrac{1}{\sqrt{2}}\left(|4p_{m},\mu\rangle\pm|4d_{m},\mu\rangle\right).$ (1)
$|\psi_{m,\mu}^{\pm}\rangle_{c}=\dfrac{1}{\sqrt{2}}\left(|4p_{m},\mu\rangle\pm i|4d_{m},\mu\rangle\right).$ (2)
These sates $|\psi_{m,\mu}^{\pm}\rangle_{p}$ and $|\psi_{m,\mu}^{\pm}\rangle_{c}$ are inspired by analogous chiral hydrogenic states [25]. Unlike hydrogen, the multielectron core potential in argon breaks inversion symmetry: the synthetic chirality is stabilized by electron correlations.
Using perturbation theory, the full spinor valued electron wave-function at the end of the ionizing pulse can be written as
$|\psi\rangle=|\psi_{o}\rangle+\sum_{I,\mu^{M}}\int d\Theta_{k}^{M}\,c_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}|I\Psi_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}^{(-)}\rangle,$ (3a)
$c_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}=i\left(\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}^{L}\cdot\boldsymbol{\vec{E}}^{L}\right)$ (3b)
where, $|\psi_{o}\rangle$ is the ground state of the molecule, $|I\Psi_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}^{(-)}\rangle$ is the fully spin-coupled continuum state with momentum $\boldsymbol{\vec{k}}^{M}$ for the ionic channel $I$, $\mu=\pm\frac{1}{2}$ labels the photoelectron spin projection onto the molecular z-axis, $\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}^{L}=\langle\Psi_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}^{(-)}|\boldsymbol{\vec{d}}^{L}|\psi_{o}\rangle$ is the spin-resolved transition dipole, and $\boldsymbol{\vec{E}}^{L}$ describes the light field [^1] Superscripts $L$ and $M$ are used to denote quantities in the laboratory and molecular frames, respectively. Vectors in the molecular frame, $\boldsymbol{\vec{a}}^{M}$, are transformed into the laboratory frame using the relation $\boldsymbol{\vec{a}}^{L}=R_{\rho}\boldsymbol{\vec{a}}^{M}$, where $R_{\rho}$ is the Euler rotation matrix. Expressions without any superscript are to be understood as fully defined in the molecular frame.. For brevity, we will drop the index $I$ in the succeeding expressions.
Figure 1: Specification of coordinates in the laboratory frame. The light field propagates along $\boldsymbol{\hat{z}}$. The unit vector $\boldsymbol{\hat{\epsilon}}$ is the polarization direction for linearly polarized light, while $\boldsymbol{\hat{\Xi}}$ is the direction of photon spin for circularly polarized light. Upon ionization, the photoelectron is ejected in the direction of $\boldsymbol{\hat{k}}$ with its spin measured parallel to $\boldsymbol{\hat{s}}$.
The momentum and spin resolved photoionization rate $W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)$ for a given orientation $\rho$ is obtained by projecting the full wavefunction Eq. (3) onto the scattering and ionic states with energy $\mathcal{E}$, and projecting the spin onto the axis $\boldsymbol{\hat{s}}^{L}$:
$\hat{\mathsf{P}}_{\mathcal{E}}$= $\sum_{\mu,\nu}\int d\Theta_{k}|\Psi_{\boldsymbol{\vec{k}},\mu}^{(-)}\rangle(\hat{\mathsf{P}}_{\boldsymbol{\hat{s}}})_{\nu,\mu}\langle\Psi_{\boldsymbol{\vec{k}},\nu}^{(-)}|.$ (4)
Here, $\hat{\mathsf{P}}_{\boldsymbol{\hat{s}}}=(\mathbb{I}+\boldsymbol{\hat{s}}\cdot\boldsymbol{\hat{\sigma}})/2$ is a spin projection operator with respect to $\boldsymbol{\hat{s}}$, and $\boldsymbol{\hat{\sigma}}$ is the vector of Pauli spin matrices. Performing the necessary operations, we obtain
$W^{L}$ $(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)$
$=$ $\dfrac{1}{2}\sum_{I,\mu,\nu}\int d\rho\left(\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\mu}^{L*}\cdot\boldsymbol{\vec{E}}^{L*}\right)\left(\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\nu}^{L}\cdot\boldsymbol{\vec{E}}^{L}\right)$
$\times\left(\delta_{\mu,\nu}+\boldsymbol{\hat{s}}^{L}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{L}\right).$ (5)
The photoelectron current conditioned on the spin detection axis $\boldsymbol{\hat{s}}^{L}$ can then be calculated as follows:
$\boldsymbol{\vec{j}}^{L}(\boldsymbol{\hat{s}}^{L})=$ $\dfrac{\int d\rho\int d\Theta_{k}^{L}W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)\boldsymbol{\vec{k}}^{L}}{\int d\rho\int d\Theta_{s}^{L}\int d\Theta_{k}^{L}W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)}$ (6)
wherein, $\int d\rho$, $\int d\Theta_{k}^{L}$, and $\int d\Theta_{s}^{L}$ denotes averaging over all molecular orientations, photoelectron momentum, and spin detection axis, respectively. Full details of the calculations are presented in the Supplementary Material.
[…]
Let us first consider randomly oriented chiral molecules under isotropic illumination by linearly polarized light:
$\boldsymbol{\vec{E}}^{L}=$ $E_{\omega}^{L}\left(\sin\theta_{p}\cos\varphi_{p}\boldsymbol{\hat{x}}^{L}+\sin\theta_{p}\sin\varphi_{p}\boldsymbol{\hat{y}}^{L}+\cos\theta_{p}\boldsymbol{\hat{z}}^{L}\right),$ (7)
where the angles $\theta_{p}$, $\varphi_{p}$ describe all possible polarization directions. Substituting Eq. (7) into (6) and averaging over all polarization directions $\int d\Theta_{p}$, the resulting photoelectron current is
$\boldsymbol{\vec{j}}_{\text{iso}}^{L}=$ $\dfrac{1}{3S_{0}}\left(\dfrac{1}{k}\int d\boldsymbol{\vec{\Theta}}_{k}^{M}\cdot\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}\right)\boldsymbol{\hat{s}}^{L},$ (8)
where,
$S_{0}=\int d\Theta_{k}\left(\left|\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},\frac{1}{2}}\right|^{2}+\left|\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},-\frac{1}{2}}\right|^{2}\right),$ (9)
is proportional to the total ionization yield. Equation (8) shows that the current is proportional to the flux of the vector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}$ through the surface of the energy shell ($d\boldsymbol{\vec{\Theta}}_{k}^{M}=d\Theta_{k}^{M}\boldsymbol{\hat{k}}^{M}k^{2}$). Mathematically, $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ is a Bloch pseudovector:
[…]
$\tilde{\varrho}=\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},\mu}^{*}\cdot\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},\nu}\right),$ (10b)
describing spin orientation in the degenerate two-level system formed by spin-up and spin-down continuum states of the photoelectron. It incorporates both populations and coherences within this system [13]. Here, $\tilde{\varrho}$ is the reduced density matrix which emerges after performing a trace over the degenerate ionization channels of the full density matrix, then averaging over all molecular orientations. Moreover, the Bloch vector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ defines a spin-texture on the energy shell for a fixed in space chiral molecule photoionized by isotropic illumination, see Fig. 2(a,b).
The current $\boldsymbol{\vec{j}}_{\text{iso}}^{L}$, Eq. (8), shows that photoelectrons with opposite spins are correlated to opposite enantiomers, i.e.,
$\boldsymbol{\vec{j}}_{\text{iso}}^{L(S)}(\boldsymbol{\hat{s}}^{L})=\boldsymbol{\vec{j}}_{\text{iso}}^{L(R)}(-\boldsymbol{\hat{s}}^{L}),$ (11)
and that the current is collinearly ‘locked’ to the direction of the spin-detection axis $\boldsymbol{\hat{s}}$: the direction of the current is collinear with spin detection axis. The strength of these correlations is defined by Eq. (8): the pseudoscalar describing spin-current correlations is a flux (or the net radial component) of the momentum-resolved Bloch vector. This expression shows the link between time-odd spin- molecular orientation correlations [13] driven by the projection of the net Bloch vector, $\boldsymbol{\vec{\mathcal{S}}}_{\text{net}}=\int d\Theta_{k}\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$, on to a molecular bond and its time-even spin-current correlations driven by the net radial component (or the flux of the momentum resolved Bloch vector through the energy shell). This result establishes momentum-resolved Bloch vector as a primary molecular property defining all pair-wise pseudoscalar correlations in photoionization of chiral molecules.
Figure 2c shows the strength of $\boldsymbol{\vec{j}}_{\text{iso}}^{L}$ for the chiral states $|\psi_{-1,\frac{1}{2}}^{\pm}\rangle_{c}$ and $|\psi_{1,\frac{1}{2}}^{\pm}\rangle_{p}$. The former shows that it can reach up to $3\%$ of the total signal while the latter shows that it is negligible. From Fig. 2(a) we see that the flux of $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ is mostly positive (negative) for the state $|\psi_{-1,\frac{1}{2}}^{+}\rangle_{c}$ ($|\psi_{-1,\frac{1}{2}}^{-}\rangle_{c}$) which is consistent with the sign of $\boldsymbol{\vec{j}}_{\text{iso}}^{L}$ in Fig. 2c. For the state $|\psi_{1,\frac{1}{2}}^{\pm}\rangle_{p}$, there is almost the same amount of positive and negative flux of the Bloch vector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ which results in a negligible $\boldsymbol{\vec{j}}_{\text{iso}}^{L}$.
The resulting photoelectron current can by enhanced by having a fixed polarization direction in the laboratory frame. Repeating the same process for linearly polarized,
$\boldsymbol{\vec{E}}^{L}=E_{\omega}^{L}\boldsymbol{\hat{\epsilon}}^{L},$ (12)
we obtain the photoelectron current
$\boldsymbol{\vec{j}}_{\text{lin}}^{L}=$ $\dfrac{1}{5S_{0}}\left[\dfrac{1}{k}\int d\boldsymbol{\vec{\Theta}}_{k}^{M}\cdot\left(2\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}+\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}\right)\right]\boldsymbol{\hat{s}}^{L}$
$-\dfrac{1}{5S_{0}}\left[\dfrac{1}{k}\int d\boldsymbol{\vec{\Theta}}_{k}^{M}\cdot\left(\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}+3\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}\right)\right]\left(\boldsymbol{\hat{s}}^{L}\cdot\boldsymbol{\hat{\epsilon}}^{L}\right)\boldsymbol{\hat{\epsilon}}^{L}$ (13a)
$\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}=$ $\text{Re}\left[\sum_{\mu,\nu}\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M*}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{M}\right)\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{M}\right].$ (13b)
The additional vector $\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}$ is the directional bias introduced by the well-defined direction of the light polarization. Hence, the photoelectron current now arises as the flux of an effective Bloch pseudovector $\boldsymbol{\vec{\mathsf{S}}}_{\boldsymbol{\vec{k}}}=a\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}+b\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}$ through the surface of the energy shell. The current $\boldsymbol{\vec{j}}_{\text{lin}}^{L}$, Eq. (13), can be measured using two possible detection geometries: (i) collinear $\boldsymbol{\hat{s}}\parallel\boldsymbol{\hat{\epsilon}}$ and (ii) orthogonal $\boldsymbol{\hat{s}}\perp\boldsymbol{\hat{\epsilon}}$, see Fig. 2c. Enhancement or reduction of $\boldsymbol{\vec{j}}_{\text{lin}}^{L}$ with respect to $\boldsymbol{\vec{j}}^{L}_{\text{iso}}$ is then controlled by $\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}$:
[…]
The comparison of the currents $\boldsymbol{\vec{j}}_{\text{iso}}^{L}$ and $\boldsymbol{\vec{j}}_{\text{lin}}^{L}$ is shown in Fig. 2c. It can be seen that the changes in the current are sensitive to both the chiral state and detection geometry.
Figure 3: Schematic of the resulting photoelectron currents for circularly polarized light in the laboratory frame. The solid and dashed blue vectors denote the opposite directions of spin for opposite enantiomers. (a,b) The current $\boldsymbol{\vec{j}}_{r}$ is enabled by the Bloch vector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ which results in a collinear locking of the photoelectron current and spin $\boldsymbol{\vec{j}}\parallel\boldsymbol{\hat{s}}$. (c,d) The currents $\boldsymbol{\vec{j}}_{\text{PECD}}$ and $\boldsymbol{\vec{j}}_{\tau}$ are enabled by the spin-resolved propensity field $\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}}$. The PECD current is not spin-sensitive, while $\boldsymbol{\vec{j}}_{\tau}$ presents a triple locking of the photoelectron current, its spin with the photon spin in three mutually orthogonal directions.
[…]
Now, consider circularly polarized light
$\boldsymbol{\vec{E}}^{L}=E_{\omega}^{L}\dfrac{(\boldsymbol{\hat{x}}^{L}+i\xi\boldsymbol{\hat{y}}^{L})}{\sqrt{2}}$ (15)
where, $\xi=\pm 1$ is a dichroic parameter characterizing the direction of rotation of the light polarization vector. The resulting current will have the following components:
$\boldsymbol{\vec{j}}_{\text{circ}}^{L}=\boldsymbol{\vec{j}}_{r}^{L}+\boldsymbol{\vec{j}}_{\text{PECD}}^{L}+\boldsymbol{\vec{j}}_{\tau}^{L}.$ (16)
The term $\boldsymbol{\vec{j}}_{r}^{L}$ is enabled by an Bloch pseudovector $\boldsymbol{\vec{\mathsf{S}}}_{\boldsymbol{\vec{k}}}$ and has a similar form as that of Eq. (13):
$\boldsymbol{\vec{j}}_{r}^{L}=$ $\dfrac{1}{10S_{0}}\left[\dfrac{1}{k}\int d\boldsymbol{\vec{\Theta}}_{k}^{M}\cdot\left(3\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}+\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}\right)\right]\boldsymbol{\hat{s}}^{L}$
$+\dfrac{1}{10S_{0}}\left[\dfrac{1}{k}\int d\boldsymbol{\vec{\Theta}}_{k}^{M}\cdot\left(\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}-3\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}\right)\right](\boldsymbol{\hat{s}}^{L}\cdot\boldsymbol{\hat{\Xi}}^{L})\boldsymbol{\hat{\Xi}}^{L},$ (17)
which is enantio-sensitive but not dichroic. The collinear locking between the photoelectron current and spin is still present and can be measured either via collinear or orthogonal detection geometry with respect to the photon spin direction $\boldsymbol{\hat{\Xi}}^{L}=(-i\boldsymbol{\vec{E}}^{L*}\times\boldsymbol{\vec{E}}^{L})/|\boldsymbol{\vec{E}}^{L}|^{2}=\xi\boldsymbol{\hat{z}}^{L}$, see Fig. 3(a,b). Moreover, the current $\boldsymbol{\vec{j}}_{r}^{L}$ is contained in the coefficients $B_{1}$ and $B_{2}$ introduced by Cherepkov [7, 8, 10]. Note that the same pseudoscalars (the fluxes of $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}$ and $\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}$) contribute with different coefficient determined by the detection geometry.
The terms $\boldsymbol{\vec{j}}_{\text{PECD}}^{L}$ and $\boldsymbol{\vec{j}}_{\tau}^{L}$ are enabled by the spin-resolved propensity field:
$(\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}})_{\mu,\nu}=i\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},\mu}^{*}\times\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},\nu}$ (18)
which is now a matrix in spin space. Explicitly, we have
[…]
which yields the familiar PECD (photoelectron circular dichroism [27, 15, 35, 17, 21, 14, 18, 19, 22, 33, 9, 28, 34, 2]) current which is not spin sensitive, presenting a “background” in a spin-conditioned measurement. Equation (19) is equal to the PECD current we derived in [24], to the original expression derived by Ritchie [30], and to the coefficient $D$ introduced by Cherepkov [7, 8, 10].
The current
$\boldsymbol{\vec{j}}_{\tau}^{L}=$ $\dfrac{1}{4N}\left(\dfrac{1}{k}\int d\boldsymbol{\vec{\Theta}}_{k}^{M}\cdot\boldsymbol{\vec{\tau}}^{M}\right)\left(\boldsymbol{\hat{s}}^{L}\times\boldsymbol{\hat{\Xi}}^{L}\right)$ (20a)
$\boldsymbol{\vec{\tau}}_{\boldsymbol{\vec{k}}^{M}}^{M}=$ $\text{Tr}(\boldsymbol{\hat{\sigma}}^{M}\times\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}^{M}}^{M}).$ (20b)
is proportional to the flux of the spin torque vector $\boldsymbol{\vec{\tau}}_{\boldsymbol{\vec{k}}^{M}}^{M}$ through the surface of the energy shell. This current is maximal when the photoelectron spin detection axis $\boldsymbol{\hat{s}}^{L}$ is orthogonal to both the direction of the current and photon spin $\boldsymbol{\hat{\Xi}}^{L}$. Thus, it is confined in the polarization plane and the respective spin-current correlations are transversal. The momentum-resolved spin-torque vector $\boldsymbol{\vec{\tau}}_{\boldsymbol{\vec{k}}^{M}}^{M}$ quantifies the intrinsic coupling between the photon and photoelectron spin mediated by the spin-resolved propensity field. $\boldsymbol{\vec{\tau}}_{\boldsymbol{\vec{k}}^{M}}^{M}$ defines the direction of the torque exerted on the photoelectron spin by the photon spin for every direction of the photoelectron momentum in the molecular frame. In this frame, the photon spin does not point along a single direction but instead forms a radially symmetric (hedgehog-like) vector field. This geometry enables access to the intrinsic photon–electron spin coupling without directional bias from the photon spin. $\boldsymbol{\vec{\tau}}_{\boldsymbol{\vec{k}}^{M}}^{M}$ exhibits skyrmionic texture (Fig. 4) which will be elucidated in our future work .
Equation (20) is also equivalent to the coefficient $C$ of Cherepkov [7, 8, 10]. The enantio-sensitive nature of $\boldsymbol{\vec{j}}_{\tau}^{L}$ arises from the molecular factor $\boldsymbol{\vec{\tau}}^{M}\cdot\boldsymbol{\vec{k}}^{M}$ (its net value, i.e. averaged over the orientation of $\boldsymbol{\vec{k}}^{M}$, see Eq.20) which is a pseudoscalar: opposite enantiomers produce opposite spin vortices, such that the spin polarization direction depends on the photoelectron momentum in a way reminiscent of Rashba effect in solids.
Alternatively, $\boldsymbol{\vec{j}}_{\tau}^{L}$ can be written in terms of the spherical multipole moments of the the spin-resolved propensity field as follows:
$\boldsymbol{\vec{j}}_{\tau}^{L}=\dfrac{1}{4S_{0}}$ $\left\{\int d\Theta_{k}^{M}\,\text{Tr}\left[\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}^{M}}^{M}\cdot\left(\boldsymbol{\vec{k}}^{M}\times\nabla_{k}\vec{\mathbb{Y}}_{\boldsymbol{\hat{k}}^{M}}^{M}\right)\right]\right\}$
$\left(\boldsymbol{\hat{s}}^{L}\times\boldsymbol{\hat{\Xi}}^{L}\right)$ (21a)
$\vec{\mathbb{Y}}_{\boldsymbol{\vec{k}}^{M}}^{M}=\sqrt{\dfrac{4\pi}{3}}\begin{bmatrix}Y_{1,0}(\boldsymbol{\hat{k}}^{M})&\sqrt{2}Y_{1,-1}(\boldsymbol{\hat{k}}^{M})\\ -\sqrt{2}Y_{1,1}(\boldsymbol{\hat{k}}^{M})&-Y_{1,0}(\boldsymbol{\hat{k}}^{M})\end{bmatrix}$ (21b)
This indicates that $\boldsymbol{\vec{j}}_{\tau}^{L}$ belongs to the Class III enantio-sensitive observables predicted in Ref. [23]. Without spin detection, multipolar enantio-sensitive currents could be induced using orthogonally polarised two-color fields [11, 31, 26, 3], and recorded via interference between the two photoionization pathways involving the interference of one-photon and two-photon transitions.
The three currents contributing to $\boldsymbol{\vec{j}}_{\text{circ}}^{L}$ (Eq.16) become mutually orthogonal when $\boldsymbol{\hat{s}}\perp\boldsymbol{\hat{\Xi}}$. In this case they can be measured independently (see Fig. 3). A comparison of each term in $\boldsymbol{\vec{j}}_{\text{circ}}^{L}$ is shown in Fig. 4c. The vortex current $\boldsymbol{\vec{j}}_{\tau}^{L}$ is negligible for the state $|\psi_{1,\frac{1}{2}}^{\pm}\rangle_{c}$ while the radial current $\boldsymbol{\vec{j}}_{r}^{L}$ is almost eight times higher than PECD. Meanwhile, $\boldsymbol{\vec{j}}_{r}^{L}$ is negligible for the state $|\psi_{-1,\frac{1}{2}}^{\pm}\rangle_{p}$ but the vortex current $\boldsymbol{\vec{j}}_{\tau}^{L}$ is almost twice as large as the PECD.
We have established two fundamental mechanisms of CISS in photoionization operating in time-even and time-odd (i.e. involving photon spin) set-ups: (i) the momentum -resolved Bloch pseudovector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ which enables collinear locking of the photoelectron current with the photoelectron spin and operates in arbitrary light fields, even under isotropic illumination, and (ii) the spin-resolved propensity field $\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}}$ which enables orthogonal triple locking between the photoelectron spin, its current and the photon spin.
We find [20] that the spin-conditioned photoelectron currents derived in Eqs. (13a), (17), and (20a) are consistent with the kinematic predictions of Cherepkov [7, 8, 10]. Our treatment makes explicit the dynamical mechanisms that were implicit in those earlier works. Importantly, it also leads to a qualitatively new prediction: the generation of a spin conditioned current even for randomly oriented chiral molecules subjected to isotropic illumination.
O.S., A.F. O. and P.C.F. acknowledge ERC-2021-AdG project ULISSES, grant agreement No 101054696. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. A.F.O. acknowledges funding from the Royal Society URF/R1/201333, URF/ERE/210358, and URF/ERE/231177 and from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 543760364.
Appendix A Derivation of photoelectron currents
The photoelectron current conditioned on the spin detection axis $\boldsymbol{\hat{s}}^{L}$ is calculated as follows:
$\boldsymbol{\vec{j}}^{L}(\boldsymbol{\hat{s}}^{L})=$ $\dfrac{\int d\rho\int d\Theta_{k}^{L}W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)\boldsymbol{\vec{k}}^{L}}{\int d\rho\int d\Theta_{s}^{L}\int d\Theta_{k}^{L}W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)}$ (22a)
$W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)=$ $\dfrac{1}{2}\sum_{\mu,\nu}\int d\rho\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{L*}\cdot\boldsymbol{\vec{E}}^{L*}\right)\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{L}\cdot\boldsymbol{\vec{E}}^{L}\right)\left(\delta_{\mu,\nu}+\boldsymbol{\hat{s}}^{L}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{L}\right)$
$=$ $\sum A^{\ell,m_{\ell}}_{\ell_{s},m_{s}}Y_{\ell,m_{\ell}}(\boldsymbol{\hat{k}}^{L})Y_{\ell_{s},m_{s}}(\boldsymbol{\hat{s}}^{L}),$ (22b)
[…]
The vectors that appear on the right-hand side of Eq. (22b) can be grouped into two sets: (i) vectors that are fixed in the molecular frame such as the dipole transition vectors $\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\mu^{M}}^{M}$, photoelectron momentum $\boldsymbol{\vec{k}}^{M}$, and photoelectron spin quantization axis $\boldsymbol{\hat{\sigma}}_{\mu_{1}^{M},\mu_{2}^{M}}^{M}$, and (ii) vectors that are fixed in the laboratory frame such as spin detection axis $\boldsymbol{\hat{s}}^{L}$ and the electric field $\boldsymbol{\vec{E}}^{L}$. This will then allow us to use the technique in Ref. [1] in evaluating the orientation averaging $\int d\rho$ such that the resulting quantity can be expressed as $\sum_{ij}g_{i}M_{ij}f_{j}$, where, $g_{i}$ and $f_{i}$ are rotational invariants that are constructed from the two sets of vectors and $M_{ij}$ is the coupling between the two rotational invariants. For our purposes, the following vector identities will be relevant:
$\int d\rho(\boldsymbol{\vec{a}}^{L}\cdot\boldsymbol{\vec{u}}^{L})\boldsymbol{\vec{b}}^{L}=\dfrac{1}{3}(\boldsymbol{\vec{a}}^{M}\cdot\boldsymbol{\vec{b}}^{M})\boldsymbol{\vec{u}}^{L}$ (23)
$\int d\rho(\boldsymbol{\vec{a}}^{L}\cdot\boldsymbol{\vec{u}}^{L})(\boldsymbol{\vec{b}}^{L}\cdot\boldsymbol{\vec{v}}^{L})\boldsymbol{\vec{c}}^{L}=\dfrac{1}{6}[(\boldsymbol{\vec{a}}^{M}\times\boldsymbol{\vec{b}}^{M})\cdot\boldsymbol{\vec{c}}^{M}](\boldsymbol{\vec{u}}^{L}\times\boldsymbol{\vec{v}}^{L})$ (24)
$\int d\rho(\boldsymbol{\vec{a}}^{L}\cdot\boldsymbol{\vec{u}}^{L})(\boldsymbol{\vec{b}}^{L}\cdot\boldsymbol{\vec{v}}^{L})(\boldsymbol{\vec{c}}^{L}\cdot\boldsymbol{\vec{w}}^{L})\boldsymbol{\vec{d}}^{L}=\dfrac{1}{30}\begin{bmatrix}(\boldsymbol{\vec{a}}^{M}\cdot\boldsymbol{\vec{b}}^{M})(\boldsymbol{\vec{c}}^{M}\cdot\boldsymbol{\vec{d}}^{M})\\ (\boldsymbol{\vec{a}}^{M}\cdot\boldsymbol{\vec{c}}^{M})(\boldsymbol{\vec{b}}^{M}\cdot\boldsymbol{\vec{d}}^{M})\\ (\boldsymbol{\vec{a}}^{M}\cdot\boldsymbol{\vec{d}}^{M})(\boldsymbol{\vec{b}}^{M}\cdot\boldsymbol{\vec{c}}^{M})\end{bmatrix}^{T}\begin{bmatrix}4&-1&-1\\ -1&4&-1\\ -1&-1&4\end{bmatrix}\begin{bmatrix}(\boldsymbol{\vec{u}}^{L}\cdot\boldsymbol{\vec{v}}^{L})\boldsymbol{\vec{w}}^{L}\\ (\boldsymbol{\vec{u}}^{L}\cdot\boldsymbol{\vec{w}}^{L})\boldsymbol{\vec{v}}^{L}\\ (\boldsymbol{\vec{v}}^{L}\cdot\boldsymbol{\vec{w}}^{L})\boldsymbol{\vec{u}}^{L}\end{bmatrix}$ (25)
Using Eqs. (23)-(24), we can easily evaluate the normalization factor as follows:
$\int d\rho$ $\int d\Theta_{s}^{L}\int d\Theta_{k}^{L}W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)$
$=$ $\int d\Theta_{s}^{L}\int d\Theta_{k}^{M}\int d\rho W^{L}(\boldsymbol{\hat{k}}^{M},\boldsymbol{\hat{s}}^{M},\rho)$
$=$ $\dfrac{1}{2}\int d\Theta_{s}^{L}\int d\Theta_{k}^{M}\sum_{\mu,\nu}\int d\rho\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{L*}\cdot\boldsymbol{\vec{E}}^{L*}\right)\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{L}\cdot\boldsymbol{\vec{E}}^{L}\right)\left(\delta_{\mu,\nu}+\boldsymbol{\hat{s}}^{L}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{L}\right)$
$=$ $\dfrac{1}{6}\int d\Theta_{s}^{L}\int d\Theta_{k}^{M}\sum_{\mu,\nu}\left[\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M*}\cdot\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{M}\right)|\boldsymbol{\vec{E}}^{L}|^{2}\delta_{\mu,\nu}\right]$
$+\dfrac{1}{12}\left\{\int d\Theta_{k}^{M}\sum_{\mu,\nu}\left[\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{M}\cdot\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M*}\times\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{M}\right)\right]\right\}\left[\left(\int d\Theta_{s}^{L}\boldsymbol{\hat{s}}^{L}\right)\cdot\left(\boldsymbol{\vec{E}}^{L*}\times\boldsymbol{\vec{E}}^{L}\right)\right]$
$=$ $\dfrac{1}{6}\left(\sum_{\mu}\int d\Theta_{k}^{M}\left|\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M}\right|^{2}\right)|\boldsymbol{\vec{E}}^{L}|^{2}$ (26)
The second line follows from the definition of a rotated function $W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)=W^{M}(\boldsymbol{\hat{k}}^{M},\boldsymbol{\hat{s}}^{M},\rho)$ [5], and we have interchanged the order of integration to perform change of variable $\boldsymbol{\hat{k}}^{M}=R_{\rho}^{-1}\boldsymbol{\hat{k}}^{L}$ and again interchanged the order of integration since $\boldsymbol{\hat{k}}^{M}$ is now an integration variable independent of $\rho$. Meanwhile, the last line follows from the vanishing of $\int d\Theta_{s}^{L}\boldsymbol{\hat{s}}^{L}$.
Let us now consider the spin-conditioned photoelectron current under isotropic illumination:
$\boldsymbol{\vec{j}}_{\text{iso}}^{L}=\dfrac{\int d\Theta_{p}\int d\rho\int d\Theta_{k}^{L}W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)\boldsymbol{\vec{k}}^{L}}{\int d\Theta_{p}\int d\rho\int d\Theta_{s}^{L}\int d\Theta_{k}^{L}W^{L}(\boldsymbol{\hat{k}}^{L},\boldsymbol{\hat{s}}^{L},\rho)},$ (27)
where, $\int d\Theta_{p}$ denotes averaging over all orientations of the light field:
$\boldsymbol{\vec{E}}^{L}_{p}=E_{\omega}^{L}\left(\sin\theta_{p}\cos\varphi_{p}\boldsymbol{\hat{x}}^{L}+\sin\theta_{p}\sin\varphi_{p}\boldsymbol{\hat{y}}^{L}+\cos\theta_{p}\boldsymbol{\hat{z}}^{L}\right)$ (28)
The relevant quantity to calculate is the the numerator of Eq. (27), i.e.,
$\int d\Theta_{k}^{M}$ $\int d\Theta_{p}\int d\rho W^{M}(\boldsymbol{\hat{k}}^{M},\boldsymbol{\hat{s}}^{M},\rho)\boldsymbol{\vec{k}}^{L}$
$=$ $\dfrac{1}{2}\sum_{\mu_{1}^{M},\mu_{2}^{M}}\int d\Theta_{k}^{M}\int d\Theta_{p}\int d\rho\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{L*}\cdot\boldsymbol{\vec{E}}_{p}^{L*}\right)\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{L}\cdot\boldsymbol{\vec{E}}_{p}^{L}\right)\left(\delta_{\mu,\nu}+\boldsymbol{\hat{s}}^{L}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}\right)\boldsymbol{\vec{k}}^{L}$ (29)
Using Eq. (24), the first term of Eq. (29) simplifies into
$\dfrac{1}{2}\sum_{\mu}$ $\int d\Theta_{k}^{M}\int d\Theta_{p}\int d\rho\left|\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{L}\cdot\boldsymbol{\vec{E}}_{p}^{L}\right|^{2}\boldsymbol{\vec{k}}^{L}$
$=$ $\dfrac{1}{12}\left\{\left[\sum_{\mu}\int d\Theta_{k}^{M}\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M*}\times\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M}\right)\right]\cdot\boldsymbol{\vec{k}}^{M}\right\}\left[\int d\Theta_{p}\left(\boldsymbol{\vec{E}}^{L*}_{p}\times\boldsymbol{\vec{E}}^{L}_{p}\right)\right]=0,$ (30)
which vanishes after averaging over all orientations of the field. Similarly, it follows from Eq. (25) that the second term of Eq. (27) is now
$\dfrac{1}{2}\sum_{\mu,\nu}\int d\Theta_{k}^{M}$ $\int d\Theta_{p}\int d\rho\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{L*}\cdot\boldsymbol{\vec{E}}_{p}^{L*}\right)\left(\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\nu}^{L}\cdot\boldsymbol{\vec{E}}_{p}^{L}\right)\left(\boldsymbol{\hat{s}}^{L}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{L}\right)\boldsymbol{\vec{k}}^{L}$
$=\dfrac{1}{60}$ $\begin{bmatrix}\sum\int d\Theta_{k}^{M}\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M*}\cdot\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{M}\right)\left(\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{M}\cdot\boldsymbol{\vec{k}}^{M}\right)\\ \sum\int d\Theta_{k}^{M}\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M*}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{M}\right)\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{M}\cdot\boldsymbol{\vec{k}}^{M}\right)\\ \sum\int d\Theta_{k}^{M}\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{M*}\cdot\boldsymbol{\vec{k}}^{M}\right)\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\nu}^{M}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{M}\right)\end{bmatrix}^{T}\begin{bmatrix}4&-1&-1\\ -1&4&-1\\ -1&-1&4\end{bmatrix}\begin{bmatrix}\int d\Theta_{p}|\boldsymbol{\vec{E}}_{p}^{L}|^{2}\boldsymbol{\hat{s}}^{L}\\ \int d\Theta_{p}(\boldsymbol{\vec{E}}_{p}^{L*}\cdot\boldsymbol{\hat{s}}^{L})\boldsymbol{\vec{E}}_{p}^{L}\\ \int d\Theta_{p}(\boldsymbol{\vec{E}}_{p}^{L}\cdot\boldsymbol{\hat{s}}^{L})\boldsymbol{\vec{E}}_{p}^{L*}\end{bmatrix}$
$=\dfrac{1}{60}$ $\begin{bmatrix}g_{1}\\ g_{2}\\ g_{2}^{*}\end{bmatrix}^{T}\begin{bmatrix}4&-1&-1\\ -1&4&-1\\ -1&-1&4\end{bmatrix}\begin{bmatrix}f_{1}\\ f_{2}\\ f_{2}^{*}\end{bmatrix}$
$=\dfrac{1}{30}$ $\left(2f_{1}-\text{Re}[f_{2}]\right)g_{1}-\dfrac{1}{30}\left(f_{1}-3\text{Re}[f_{2}]\right)\text{Re}[g_{2}]-\dfrac{1}{6}\text{Im}[f_{2}]\text{Im}[g_{2}]$ (31)
Averaging over all orientations of the field we get:
$\dfrac{1}{2}\sum_{\mu,\nu}\int d\Theta_{k}^{M}$ $\int d\Theta_{p}\int d\rho\left(\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}^{M},\mu}^{L*}\cdot\boldsymbol{\vec{E}}_{p}^{L*}\right)\left(\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\nu}^{L}\cdot\boldsymbol{\vec{E}}_{p}^{L}\right)\left(\boldsymbol{\hat{s}}^{L}\cdot\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{L}\right)\boldsymbol{\vec{k}}^{L}$
$=$ $\dfrac{1}{18}|\boldsymbol{\vec{E}}^{L}|^{2}\left\{\int d\Theta_{k}^{M}\left[\sum_{\mu,\nu}\left(\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\mu}^{M*}\cdot\boldsymbol{\vec{D}}_{I,\boldsymbol{\vec{k}}^{M},\nu}^{M}\right)\boldsymbol{\hat{\sigma}}_{\nu,\mu}^{M}\right]\cdot\boldsymbol{\vec{k}}^{M}\right\}\boldsymbol{\hat{s}}^{L}$
$=$ $\dfrac{1}{18}|\boldsymbol{\vec{E}}^{L}|^{2}\left[\int d\Theta_{k}^{M}\left(\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}\cdot\boldsymbol{\vec{k}}^{M}\right)\right]\boldsymbol{\hat{s}}^{L}.$ (32)
Thus,
$\boldsymbol{\vec{j}}_{\text{iso}}^{L}=\dfrac{1}{3S_{0}}\left[\int d\Theta_{k}^{M}\left(\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}\cdot\boldsymbol{\vec{k}}^{M}\right)\right]\boldsymbol{\hat{s}}^{L}$ (33a)
$S_{0}=\int d\Theta_{k}\left(\left|\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},\frac{1}{2}}\right|^{2}+\left|\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},-\frac{1}{2}}\right|^{2}\right).$ (33b)
The same procedure is done for both linearly and circularly polarized light.